A sound wave travels from bone(Z =8.0 x 10^6 kgm-2s-1)to water . What will the intensity of the reflected wave be as a percentage of the intensity of the incident waves?


1
Expert's answer
2021-03-26T11:40:43-0400

Reflection fraction


"\frac{(Z_2-Z_1)^2}{(Z_2+Z_1)^2}=\frac{(8-1.48)^2}{(8+1.48)^2}=0.473=47.3\%"